# A Comprehensive Evaluation of Parameter-Efficient Fine-Tuning on Method-Level Code Smell Detection

### v1

@@ 5. Discussion
Approaches with improved effectiveness for CC code smell detection are needed: According to Fowler (fowler1999refactoring, ), four factors contribute to making programs difficult to work with, one of which is Complex Conditional (CC) logic that complicates modification. The presence of CC in source code hinders testing and maintenance (fowler1999refactoring, ; liu2014automated, ; alhefdhi2022framework, ). Thus, identifying CC code smells in software can significantly improve code quality (liu2014automated, ). However, the effectiveness of existing methods for detecting this type of code smell is considerably limited. As shown in Table 8, existing methods — whether heuristic-based (e.g., DesigniteJava), DL-based (e.g., DeepSmells and AE-Dense), or GPT-3.5-turbo — demonstrate limited success in detecting CC, with all MCC values below 45%, and some even as low as 6.32%. Although our proposed method, GraphCodeBERT with $(IA)^{3}$, improved the performance of CC code smell detection significantly, reaching an F1-score of 77.58% and an MCC value of 56.91%, it still falls short of expectations due to challenges such as the considerable existence of false positives and negatives, which may limit its practical applicability. Accurate identification of CC code smells can help developers better refactor the code, ultimately improving the maintainability of the software system. Despite the progress made, the detection of CC remains a challenging task that requires further improvements.
Model size is not the decisive factor contributing to the effectiveness of code smell detection: Previous research has demonstrated that larger models with more parameters tend to deliver better performance on code-related tasks (nijkamp2022codegen, ; roziere2024codellama, ; weyssow2023exploring, ). For example, Weyssow et al. (weyssow2023exploring, ) suggested a potential shift in the software engineering domain, moving away from the long-established practice of fine-tuning smaller LMs like CodeT5 over the past 3-4 years, and towards embracing larger LLMs. However, the results of RQ1 indicate that, for the method-level code smell detection task, model size is not the sole determinant of performance. In fact, all the small models, with the exception of CodeT5, generally perform as well as or better than larger LLMs in detecting CC and CM (see Section 4.1.3). Furthermore, within the same model family, larger models do not always outperform its smaller-sized counterparts (see Section 4.1.2). In contrast, the alignment between the model and the specific task at hand seems to play a crucial role in determining effectiveness for code smell detection. While CodeT5, when fine-tuned with PEFT methods, has demonstrated strong performance in tasks such as code summarization and code change prediction (liu2023empirical, ; fan2024exploring, ), it surprisingly underperforms in the task of code smell detection. This suggests that models may exhibit varying performance across different downstream tasks, emphasizing the importance of aligning a model’s design and fine-tuning with the specific requirements of each target task. Therefore, it is essential to focus on selecting models that are well-suited to the method-level code smell detection task, rather than relying solely on model size as a performance indicator.
Considering the model used and the availability of resources when selecting the appropriate PEFT method for code smell detection: Based on the results of RQ1 and RQ3, we can find that for method-level code smell detection, the most effective PEFT method can differ significantly depending on the models and the amount of available data. From Table 5 and Table 6, it is obvious that each model responds to different PEFT techniques in unique ways, suggesting that the optimal PEFT method is rather model-dependent. The effectiveness of PEFT methods can change even for the same model under different resource scenarios. For instance, in low-resource settings with limited data, the best PEFT method for models like GraphCodeBERT and DeepSeek-Coder-1.3B shifts from prefix tuning to LoRA. Moreover, peak GPU consumption can vary across different models (see Section 4.1.2). Generally, $(IA)^{3}$ requires more GPU resources than LoRA for code smell detection, whereas the reverse is true for StarCoderBase-7B. These variations underscore that there is no universally best PEFT method for method-level code smell detection; instead, the most suitable method depends on the specific model and resource constraints. Therefore, we recommend that practitioners consider both the model’s characteristics and the availability of resources, including data volume and GPU memory, when selecting the appropriate PEFT method for a given task.
@@
Start model selection for code smell detection with a smaller dataset to optimize resource usage: In low-resource scenarios, when the dataset size was reduced by more than 90%, the models’ effectiveness did not degrade significantly, and they still yielded promising results, as shown in the results of RQ3 (see Section 4.3). This finding indicates that, even with a significantly smaller dataset, both small and large LMs can maintain robust performance for the code smell detection task. Consequently, beginning fine-tuning experiments with a smaller-size dataset can be an effective strategy to optimize resource usage. This approach allows researchers and practitioners to initially evaluate a wide range of LMs in a shorter time and wither fewer computational resources, saving substantial effort. It enables them to identify the most suitable model for the specific task through an initial evaluation before committing to the more resource-intensive process of training on larger datasets. Additionally, it helps in assessing the trade-offs between resource consumption and performance, providing insight into the models’ scalability. By leveraging smaller datasets for initial evaluation, one can make informed decisions about which models are likely to perform well with limited resources, and further fine-tune or scale up as necessary, without wasting valuable time or computational power.
@@ 6. Threats to Validity
First, relying solely on DesigniteJava for initial sample detection may introduce bias. However, DesigniteJava is a well-established and widely adopted tool in Java code smell detection (chaniotaki2021architecture, ; sharma2021qscored, ; shah2023mining, ), providing a robust foundation for the identification of potential code smells. We also acknowledge that future studies could benefit from incorporating additional tools in the initial code smell detection process, which may help mitigate the threat to construct validity by providing more diverse perspectives on the data.
Second, the manual validation process, while aimed at ensuring dataset accuracy, could introduce some subjectivity due to its reliance on human judgment. To minimize this threat, we conducted separate pilot validations for CC and CM before formally annotating all the data samples. This pilot phase helped ensure that the two annotators used the same criteria for validation, thus improving consistency and reducing potential bias in the dataset labeling.
External Validity: In this study, we evaluated the effectiveness of state-of-the-art PEFT methods using 4 small LMs and 6 LLMs. However, due to limited computational resources, we did not include larger models with over 7B parameters, which may impact the generalizability of our findings. Additionally, we only employed four popular and typical PEFT techniques, i.e., prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$. Other PEFT methods, such as adapter tuning (houlsby2019parameter, ), could be explored in our future work. Moreover, our study only applied PEFT methods to detect Java code smells. Expanding this research to include code smells in other programming languages, such as C# and Python, could provide further insights into the effectiveness and adaptability of PEFT methods across diverse language-specific code smell patterns.
@@ 7.1. Code Smell Detection
Traditional heuristic-based code smell detectors face inherent challenges, such as difficulty in defining effective thresholds for detection. In response, Machine Learning (ML) and Deep Learning (DL) techniques have gained traction as promising alternatives. Maiga et al. (maiga2012support, ) implemented SVMDetect based on an ML technique — support vector machines. Fontana et al. (arcelli2016comparing, ) evaluated 16 ML algorithms across four types of code smells using 74 software systems, reporting that J48 and Random Forest delivered the highest performance.
Liu et al. (liu2021deep, ) utilized deep neural networks and advanced DL techniques to perform useful feature selection from source code and map these features to labels (smelly or not). Sharma et al. (sharma2021code, ) trained Convolution Neural Network (CNN), Recurrent Neural Network (RNN), and autoencoder DL models in different configurations to detect code smells. Their work confirms the feasibility of applying transfer-learning for code smell detection. Ho et al. (ho2023fusion, ) developed DeepSmells based on a CNN and long short-term memory (LSTM) networks for detecting code smells. They conducted an empirical study to compare DeepSmells with other well-established detection tools. Their results demonstrate that DeepSmells outperforms the baselines, achieving superior performance. They also anticipated that leveraging CodeBERT would enhance the effectiveness of code smell detection. Recently, Wu et al. (wu2024ismell, ) trained a Mixture of Experts (MoE) model, leveraging specialized expert toolsets for code smell detection. They subsequently utilized the detection results to prompt LLMs for refactoring.
Most prevailing ML- or DL-based approaches for code smell detection train models from scratch. Compared to their works, we harness the power of LLMs using PEFT methods to detect two types of complex method-level code smells (Complex Conditional and Complex Method), aiming to enhance the models’ ability to capture the nuanced relationships between different types of method-level code smells.
@@ 7.2. PEFT Methods on Software Engineering Tasks
Several studies have evaluated PEFT methods on various SE tasks. Wang et al. (wang2023one, ) fine-tuned UniXcoder and CodeT5 by inserting the parameter-efficient structure adapter for code search and summarization. According to their results, adapter tuning has shown effectiveness in cross-lingual and low-resource scenarios, surpassing full fine-tuning and effectively mitigating catastrophic forgetting. Liu et al. (liu2023empirical, ) carried out experiments by fine-tuning code models with adapter, prefix tuning, LoRA, and MHM (he2022towards, ) on both the code understanding task and code generation task. Their findings indicate that adapter tuning and LoRA can achieve promising performance on these two SE tasks with reduced training costs and memory requirements. Liu et al. (liu2024delving, ) examined whether adapter tuning and LoRA can perform well on code-change-related tasks, specially just-in-time defect prediction and commit message generation. Their study demonstrates that PEFT methods can serve as a powerful alternative to full fine-tuning for dynamic code comprehension. Weyssow et al. (weyssow2023exploring, ) compared PEFT methods, including prompt tuning and others, with In-Context Learning (ICL) in terms of code generation. Their investigation reveals the superiority of PEFT over ICL and highlights the potential opportunities of tuning LLMs in diverse software engineering scenarios. Li et al. (li2024comprehensive, ) created an instruction dataset of Automated Program Repair (APR) using prompt engineering. They employed this dataset to fine-tune four LLMs using four different PEFT methods. The results show that $(IA)^{3}$ achieves the highest fixing ability compared to prompt tuning, prefix tuning, and LoRA.
These studies focus on evaluating PEFT methods across a range of software engineering tasks, such as code summarization and code generation. Compared to their works, our work pays specific attention to a binary classification task, aiming to assess the effectiveness of LLMs with PEFT methods in identifying different types of code smells.

### v2

@@ 5. Discussion
Approaches with improved effectiveness for CC code smell detection are needed: According to Fowler (fowler1999refactoring, ), four factors contribute to making programs difficult to work with, one of which is Complex Conditional (CC) logic that complicates modification. The presence of CC in source code hinders testing and maintenance (fowler1999refactoring, ; liu2014automated, ; alhefdhi2022framework, ). Thus, identifying CC code smells in software can significantly improve code quality (liu2014automated, ). However, the effectiveness of existing methods for detecting this type of code smell is considerably limited. As shown in Table 10, existing methods, including heuristic-based (e.g., DesigniteJava), DL-based (e.g., DeepSmells) and AE-Dense), and general-purpose LLMs under ICL settings (e.g., GPT-4o-mini and DeepSeek-v3), all demonstrate limited success in detecting CC, with all MCC values below 50%, and some even as low as 6.63%. Although our proposed method, StarCoderBase-3B with $(IA)^{3}$, improved the performance of CC code smell detection significantly, reaching an F1-score of 74.68% and a MCC value of 49.96%, it still falls short of expectations due to challenges such as the considerable existence of false positives and negatives, which may limit its practical applicability. Accurate identification of CC code smells can help developers better refactor the code, ultimately improving the maintainability of the software system. Despite the progress made, the detection of CC remains a challenging task that requires further improvements.
PEFT methods can enhance the capacity of LMs in class-level code understanding: In this study, we selected four representative types of code smells, i.e., CC, CM, FE, and DC, to evaluate the effectiveness of PEFT methods. While only DC is formally categorized as a class-level smell, FE detection also inherently requires class-level analysis as well, as it involves analyzing method-class dependencies (see Section 3.3.1 and Section 3.3.4). According to the recent studies (silva2024detecting, ; mesbah2025leveraging, ), LLMs not specifically fine-tuned for code smell detection generally perform worse when tasked with identifying smells that demand class-level understanding. This aligns with our intuitive expectation that class-level code understanding is more difficult, as it requires capturing higher-level structural relationships (e.g., field-method interactions, class cohesion, and design-level abstractions). Despite this challenge, most existing research on PEFT methods has predominantly concentrated on the their perfomance for method-level code understanding (liu2023empirical, ; wang2023one, ; liu2024delving, ), leaving their effectiveness on class-level tasks largely unexplored. Our work bridges this critical gap by evaluating PEFT methods on code smell detection, a task spans both method- and class-level analysis. Surprisingly, as discussed in Section 4.1.5, our results reveal that LLMs with PEFT methods not only overcome the anticipated class-level challenges but even outperform method-level detection in some cases. These findings suggest that PEFT methods not only improve parameter efficiency but also strengthen the proficiency of LLMs to perform robust code understanding across different levels of granularity.
While larger LMs with PEFT methods often achieve better effectiveness in code smell detection, SLMs still retain distinct advantages: The results of RQ1 reveal that LLMs generally outperform SLMs in code smell detection, particularly in identifying FE. This performance advantage aligns with established neural scaling laws (kaplan2020scaling, ), which posit that model capabilities improve predictably with increased scale. However, SLMs are able to achieve comparable performance in detecting CC and CM, and even significantly outperform LLMs in the detection of DC. Moreover, as illustrated in Figure 5, SLMs exhibit substantially lower peak GPU memory consumption compared to LLMs, making them a more resource-efficient choice. These findings highlight a critical trade-off: while LLMs tend to yield higher absolute performance, SLMs offer a competitive and cost-effective alternative, especially in settings with limited computational resources.
Considering the model used, the code smell type, and the availability of resources when selecting the appropriate PEFT method: Based on the results of RQ1 and RQ3, we can find that for code smell detection, the most effective PEFT method can differ significantly depending on the models and the amount of available data. From Table 6 and Table 7, we observe that although $(IA)^{3}$ generally achieves the best performance for LLMs across most code smell types, this is not universally the case. For certain models, particularly SLMs, and specific types of code smells, prefix tuning or even prompt tuning may outperform $(IA)^{3}$. As indicated in Figure 5, the relatively high peak GPU memory consumption of $(IA)^{3}$ may limit its practicality in constrained environments, where more lightweight PEFT methods could be more suitable. Moreover, in low-resource scenarios where training data is limited, LoRA emerges as the most effective choice (see results of RQ3 in Section 4.3.2). These findings emphasize the importance of selecting PEFT strategies holistically, taking into account not only the model size and available resources, including both GPU memory and training data, but also the specific demands of the code smell being detected.
@@
Start model selection for code smell detection with a smaller dataset to optimize resource usage: In low-resource scenarios, even when the dataset size was drastically reduced, the models’ effectiveness did not degrade significantly, and they still yielded promising results, as shown in the results of RQ3. This finding indicates that, even with a significantly smaller dataset, both small and large LMs can maintain robust performance for the code smell detection task. Consequently, beginning fine-tuning experiments with a smaller-size dataset can be an effective strategy to optimize resource usage. This approach allows researchers and practitioners to initially evaluate a wide range of LMs in a shorter time and wither fewer computational resources, saving substantial effort. It enables them to identify the most suitable model for the specific task through an initial evaluation before committing to the more resource-intensive process of training on larger datasets. Additionally, it helps in assessing the trade-offs between resource consumption and performance, providing insight into the models’ scalability. By leveraging smaller datasets for initial evaluation, one can make informed decisions about which models are likely to perform well with limited resources, and further fine-tune or scale up as necessary, without wasting valuable time or computational power.
ICL shows promise as an effective alternative to PEFT methods for certain types of code smells: In RQ4, we investigate the performance of general-purpose LLMs in code smell detection under In-Context Learning (ICL) settings. As shown in Table 10, for CC and DC, the optimal ICL configuration of DeepSeek-v3 achieves MCC scores only 4%–5% lower than those attained by the top-performing PEFT-tuned LMs. This suggests that, for certain types of code smells, well-crafted prompts and a few in-context examples may enable general-purpose LLMs to produce competitive results. Moreover, applying ICL technique requires no model training or parameter updates (zhou2024large, ; tom2020language, ; li2025large, ), which is particularly appealing for resource-constrained scenarios or when model weights cannot be modified (e.g., in commercial LLM APIs). However, for CM and FE, ICL lags significantly behind, with performance deficits of up to 13.69% and 29.25% in MCC, respectively, compared to our PEFT-based approaches. These results indicate that detecting these smell types likely requires deeper model adaptation to capture subtle semantic patterns and structural nuances that few-shot prompting alone cannot convey. Overall, while ICL offers a lightweight and flexible alternative, especially for rapid prototyping or scenarios with limited computational resources, PEFT techniques remain the more robust and reliable solution for achieving high-accuracy code smell detection across a broad range of smell types.
Future work could explore the use of Retrieval-Augmented Generation (RAG) for code smell detection: According to Figure 7, while models generally perform better in code smell detection when provided with randomly selected examples compared to the zero-shot setting, their performance does not consistently improve as the number of example pairs increases. This observation suggests that simply increasing the number of in-context examples without considering their quality, relevance, or diversity may not be sufficient to enhance model effectiveness for the task of code smell detection. By contrast, Retrieval-Augmented Generation (RAG) (lewis2020retrieval, ) can dynamically retrieve the most relevant examples from an external knowledge base to augment the prompts. By grounding predictions in high-quality and contextually appropriate evidence, RAG has the potential to enhance detection accuracy while avoiding the need for fine-tuning or complex prompt engineering. As such, future research could explore whether RAG-based approaches can provide a more effective and scalable solution to PEFT methods for code smell detection.
@@ 6. Threats to Validity
First, relying solely on DesigniteJava and PMD for initial sample detection may introduce bias. However, DesigniteJava and PMD are well-established and widely adopted tools in Java code smell detection (chaniotaki2021architecture, ; sharma2021qscored, ; shah2023mining, ; tushar2016does, ; taibi2017how, ), providing a robust foundation for the identification of potential code smells. We also acknowledge that future studies could benefit from incorporating additional tools in the initial code smell detection process, which may help mitigate the threat to construct validity by providing more diverse perspectives on the data.
Second, the manual validation process, while aimed at ensuring dataset accuracy, could introduce some subjectivity due to its reliance on human judgment. To minimize this threat, we conducted separate pilot validations for the four types of code smells before formally annotating all the data samples. During this phase, we established and refined clear identification criteria to ensure alignment between annotators. This approach not only improved labeling consistency but also reduced potential bias in the dataset.
Lastly, due to budget constraints for commercial API usage, our evaluation of general-purpose LLMs (GPT-4o-mini and DeepSeek-v3) under ICL settings in RQ4 was limited to a maximum of three example pairs. This sample size may not fully represent the upper bounds of ICL performance. Future work could expand on this by increasing the number of example pairs and evaluating additional models to better understand the full potential of ICL performance. However, prior studies such as UniLog (xu2024unilog, ) have shown that the performance of LLMs on software engineering tasks does not always improve linearly with more examples, which suggests that while increasing the number of examples may be beneficial, performance gains are not guaranteed. As such, we believe our current evaluation still provides valuable insights and that this threat to validity is partially mitigated.
External Validity: In this study, we evaluated the effectiveness of state-of-the-art PEFT methods using 4 small LMs and 5 LLMs. However, due to limited computational resources, we did not include larger models with over 7B parameters, which may impact the generalizability of our findings.
Furthermore, we only employed four widely used and typical PEFT techniques, i.e., prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$. While these methods cover the major categories defined in the PEFT library, the exclusion of other approaches may limit the generalizability of our findings. For example, methods such as adapter tuning (houlsby2019parameter, ) could be explored in future work to provide a more comprehensive evaluation.
Moreover, our study applied PEFT methods exclusively to the detection of Java code smells, focusing on four specific types: three method-level smells (Complex Conditional, Complex Method, and Feature Envy) and one class-level smell (Data Class). While this selection provides a certain degree of diversity, it does not fully capture the wide spectrum of code smell types encountered in real-world software systems. Expanding future work to include a wider range of smell categories would enhance the comprehensiveness of the evaluation. In addition, our current work is limited to Java. Applying PEFT techniques to detect code smells in other programming languages, such as C# or Python, could yield further insights into their generalizability and adaptability across different language ecosystems.
@@
Another potential threat to internal validity stems from conducting only a single experimental run per configuration due to computational resource constraints. Although we ensured consistency across all experiments by using a fixed random seed (42), this approach cannot fully account for the inherent stochasticity in the training process. To assess the impact of randomness, we conducted supplementary experiments using two alternative seeds (0 and 999) for a representative case (GraphCodeBERT on CC detection). The results shown in Table 11 indicate that there are only small variations in the performance metrics across different random seeds. These minor differences suggest that our core findings are robust to random initialization effects, with the overall trends remaining consistent. While some performance fluctuations are observed, they do not significantly alter the conclusions drawn from the experiments. However, we acknowledge that conducting multiple runs with different seeds would yield more statistically reliable conclusions.
Table 11. Results of GraphCodeBERT with PEFT methods and full fine-tuning on detecting CC across different seeds
Seed Method Precision Recall F1 MCC
0 Full FT 73.47% 72.17% 71.79% 45.62%
Prompt Tuning 64.89% 64.89% 64.89% 29.78%
Prefix Tuning 71.32% 71.30% 71.30% 42.62%
LoRA 73.17% 73.04% 73.01% 46.21%
$(IA)^{3}$ 72.44% 72.17% 72.09% 44.61%
42 Full FT 74.06% 74.02% 74.01% 48.08%
Prompt Tuning 65.49% 65.33% 65.24% 30.81%
Prefix Tuning 70.08% 69.89% 69.82% 39.97%
LoRA 73.15% 73.15% 73.15% 46.30%
$(IA)^{3}$ 72.50% 72.50% 72.50% 45.00%
999 Full FT 73.19% 72.83% 72.72% 46.02%
Prompt Tuning 64.26% 64.24% 64.22% 28.50%
Prefix Tuning 71.21% 69.57% 68.96% 40.74%
LoRA 73.07% 72.93% 72.90% 46.00%
$(IA)^{3}$ 72.23% 72.17% 72.15% 44.41%
@@ 7.1. Code Smell Detection
Traditional heuristic-based code smell detectors face inherent challenges, such as difficulty in defining effective thresholds for detection. In response, Machine Learning (ML) and Deep Learning (DL) techniques have gained traction as promising alternatives. Maiga et al. (maiga2012support, ) implemented SVMDetect based on an ML technique — support vector machines. Fontana et al. (fontana2016comparing, ) evaluated 16 ML algorithms across four types of code smells using 74 software systems, reporting that J48 and Random Forest delivered the highest performance.
Liu et al. (liu2021deep, ) utilized deep neural networks and advanced DL techniques to perform useful feature selection from source code and map these features to labels (smelly or not). Sharma et al. (sharma2021code, ) trained Convolution Neural Network (CNN), Recurrent Neural Network (RNN), and autoencoder DL models in different configurations to detect code smells. Their work confirms the feasibility of applying transfer-learning for code smell detection. Ho et al. (ho2023fusion, ) developed DeepSmells based on a CNN and long short-term memory (LSTM) networks for detecting code smells. They conducted an empirical study to compare DeepSmells with other well-established detection tools. Their results demonstrate that DeepSmells outperforms the baselines, achieving superior performance. They also anticipated that leveraging CodeBERT would enhance the effectiveness of code smell detection.
We conduct an extensive analysis of the Machine Learning Code Quality (MLCQ) dataset, focusing on how these LLMs perform when prompted to identify and classify code smells.
Recently, researchers have started to explore the use of LLMs for code smell detection. Mesbah et al. (mesbah2025leveraging, ) used the MLCQ dataset to analyze how GPT-4 and LLaMA perform when prompted to identify and classify code smells. Silva et al. (silva2024detecting, ) employed two types of prompts to query ChatGPT for identifying Blob, Data Class, Feature Envy, and Long Method code smells. Sadik et al. (sadik2025benchmarking, ) created a dataset for evaluating LLMs on automated code smell detection. They benchmarked two state-of-the-art LLMs and found that LLMs are not yet ready to replace deterministic tools in all aspects of code quality assurance. Wu et al. (wu2024ismell, ) trained a Mixture of Experts (MoE) model, leveraging specialized expert toolsets for code smell detection. They subsequently utilized the detection results to prompt LLMs for refactoring.
Most prevailing ML-based or DL-based approaches for code smell detection train models from scratch. On the other hand, existing LLM-based approaches typically rely on simple prompting strategies without any form of fine-tuning or adaptation. Compared to their works, we harness the power of LLMs using PEFT methods to detect four types of widely studied code smells (Complex Conditional, Complex Method, Feature Envy, and Data Class), aiming to enhance the models’ ability to capture the nuanced relationships between different code smell types.
@@ 7.2. PEFT Methods on Software Engineering Tasks
Several studies have evaluated PEFT methods on various Software Engineering (SE) tasks. Wang et al. (wang2023one, ) fine-tuned UniXcoder and CodeT5 by inserting the parameter-efficient structure adapter for code search and summarization. According to their results, adapter tuning has shown effectiveness in cross-lingual and low-resource scenarios, surpassing full fine-tuning and effectively mitigating catastrophic forgetting. Liu et al. (liu2023empirical, ) carried out experiments by fine-tuning code models with adapter, prefix tuning, LoRA, and MHM (he2022towards, ) on both the code understanding task and code generation task. Their findings indicate that adapter tuning and LoRA can achieve promising performance on these two SE tasks with reduced training costs and memory requirements. Liu et al. (liu2024delving, ) examined whether adapter tuning and LoRA can perform well on code-change-related tasks, specially just-in-time defect prediction and commit message generation. Their study demonstrates that PEFT methods can serve as a powerful alternative to full fine-tuning for dynamic code comprehension. Weyssow et al. (weyssow2023exploring, ) compared PEFT methods, including prompt tuning and others, with In-Context Learning (ICL) in terms of code generation. Their investigation reveals the superiority of PEFT over ICL and highlights the potential opportunities of tuning LLMs in diverse SE scenarios. Li et al. (li2024comprehensive, ) created an instruction dataset of Automated Program Repair (APR) using prompt engineering. They employed this dataset to fine-tune four LLMs using four different PEFT methods. The results show that $(IA)^{3}$ achieves the highest fixing ability compared to prompt tuning, prefix tuning, and LoRA.
These studies focus on evaluating PEFT methods across a range of SE tasks, primarily at the method level. Compared to their works, our work specifically targets the task of code smell detection, which, to the best of our knowledge, has not been studied before. Additionally, our work further explores the capacity of LLMs on class-level code understanding, thereby addressing a significant gap in existing research.